OpenStudy (anonymous):

Coin can be tossed until a tail appears or until it has been tossed 3 times. Given that tail does not occur on the first toss, whats probability that the coin is tossed 3 times? I think I got answer by thinking but i don't know how to write it out So I need to use conditional probability and P(no tail on 1 try)=0.5 and how to get others?

OpenStudy (anonymous):

you already know that you didn't have a tail on the first toss, so change the words "3 times" to "two times"

OpenStudy (anonymous):

sample space is now HHH HHT HT

OpenStudy (anonymous):

Nice @ satellite 73

OpenStudy (anonymous):

in two out of the 3 cases it is thrown three times, unless i misinterpreted the question

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

given answer is actually 1/2

OpenStudy (anonymous):

if you want a formula, you can write A= tail is tossed on first tryt B= coin is tossed 3 times and then you want \[P(B|A)=\frac{P(A\cap B)}{P(A)}\] but this begs the question because you still have to compute \[P(A\cap B)\] which just requires the same work, computing \[A\cap B=\{HHH, HHT\}\]

OpenStudy (anonymous):

i had a typo above, it should have been A = heads tossed on first try ok first of all the answer, and it is \[\frac{2}{3}\] because out of the three events HHH HHT HT two are favorable

OpenStudy (anonymous):

no \[P(A\cap B)\neq \frac{1}{2}\]

OpenStudy (anonymous):

experiment is "Coin can be tossed until a tail appears or until it has been tossed 3 times." so lets write out a complete sample space, as it is small

OpenStudy (anonymous):

shouldn't we write complete sample T, HT, HHT, HHH and given answer is 1/2 not 2/3 but it might be incorrect

OpenStudy (anonymous):

{T, HT, HHT, HHH} only 4 possiblities

OpenStudy (anonymous):

exactly what you wrote yes

OpenStudy (anonymous):

so if we put the event A = first toss is heads, we know \[P(A)=\frac{1}{2}\]easily, but what we need is \[P(B|A)=\frac{P(A\cap B)}{P(A)}\]

OpenStudy (anonymous):

\[\frac{3}{4}\div \frac{1}{2}>1\]

OpenStudy (anonymous):

well i made a mistake of course, (not in the answer, in the computation)

OpenStudy (anonymous):

P(A) is correct, how to get P(A∩B)?

OpenStudy (anonymous):

if we look at full sample T, HT, HHT, HHH then HHT and HHH suits A and B, no? but it means 0.5 and 0.5/0.5=1 and again it's incorrect :/

OpenStudy (anonymous):

ok i see the problem

OpenStudy (anonymous):

i am sorry i am an idiot this morning, so lets go slow

OpenStudy (anonymous):

first of all there are 4 possible outcomes, T HT HHT HHH but they are not EQUALLY LIKELY so we cannot use the uniform distribution

OpenStudy (anonymous):

that is, each does not occur with probability 1/4

OpenStudy (anonymous):

\[P(T)=\frac{1}{2}, P(HT)=\frac{1}{4},P(HHH)=\frac{1}{8},P(HHH)=\frac{1}{8}\]

OpenStudy (anonymous):

ah, yeah..

OpenStudy (anonymous):

ok now i understood lol

OpenStudy (anonymous):

so i was completely wrong, and i apologize. should have a nice cup of coffee

OpenStudy (anonymous):

ok so now let us finish the problem.

OpenStudy (anonymous):

so \[ P(A \cap B)=\frac{1}{8}+\frac{1}{8}=\frac{1}{4}\]

OpenStudy (phi):